how do you get the answer to (-3+2i)^2
Found 2 solutions by flame8855, lynnlo:
Answer by flame8855(424)

= 9 +4i^2 -12i
9-4-12i
5-12i

Answer by lynnlo(4176)

(-3+2i)^2=5-12i